OpenStudy (anonymous):

2x+20=x^2+10x .. solve for x.... PLEASE HELPP

OpenStudy (anonymous):

\[2x+20=x^2+10x\] Move everything to the LHS (Left Hand Side) \[x^2 +8x -20=0\] Factorise the quadratic equation. \[(x+10)(x-2)=0\] (therefore) x=-10 or x=2

OpenStudy (anonymous):

Thank you!